# supplemental

### v1

@@ Abstract
with all the coefficients $a_{i}=a_{i}(\lambda,M)$. Due to the system not possessing an $R\to-R$ symmetry, we have to include all terms of the expansion. Related Landau-like expansions have been made in [24, 29]. The expansion of $\langle C\rangle$ is obtained directly from (3), (4) and (6), (7). We also require the expansion of $P_{\rm s}$ for the two resetting ensembles, which is obtained by expanding out the integrands of (5) and (8) in terms of $R$ and integrating term by term. This procedure is carried out in detail in the SM [35].
@@
The simplest case is when $a_{2}>0$ in the expansion (9) and we may ignore higher order terms. The curve of continuous transitions, $\lambda^{*}(M)$, is obtained by solving $a_{1}(\lambda,M)=0$ for $\lambda$. This is the usual scenario for a continuous transition seen in equilibrium systems [41]. This simple scenario pertains in the case of RBM with a linear cost, where we saw in Fig. 2 that a continuous transition occurs on increasing $\lambda$, for sufficiently low $M$. The threshold value $M_{\rm T}$ is given by $\lambda^{*}(M_{\rm T})\to\infty$ and for $M>M_{T}$ no transition occurs. The exact value $M_{\rm T}=0.8198\ldots$ that is obtained from this procedure provides the horizontal line in Fig. 2(a).
@@
The next case is where $a_{2}$ may be positive or negative according to parameters, but $a_{3}$ is positive. This is the case for RBB with linear cost, Fig. 2(c). Then for $a_{2}>0$ we obtain a continuous transition at $a_{1}=0$ but for $a_{2}<0$ there is a discontinuous transition. A classical tricritical point occurs when $a_{1}=a_{2}=0$ and $a_{3}>0$. To obtain the tricritical point, we set $a_{2}(\lambda^{*},M)=0$, which yields $(M^{*},\lambda^{*})=(0.6333,8.4994)$ indicated by the filled circle in Fig. 2(c).
@@
The interesting non-classical overhangs that occur for both RBM and RBB with quadratic cost per reset (Figs. 2(b) and 2(d)) can be attributed to the coefficients satisfying $a_{3}<0$ and $a_{4}>0$. If we look for a tricritical point by solving $a_{2}(\lambda^{*},M)=0$ for $M$, we find at all such $M^{*}$ that $a_{3}(\lambda^{*},M^{*})<0$. This violates the condition for a tricritical point, which is why overhangs emerge instead. Of course the predictions of a Landau-like theory for discontinuous transitions will only be quantitatively accurate for small $R$, nevertheless the correct qualitative behaviour is predicted.
@@
In summary, we have introduced the profligacy $\xi$ (1) as a tool to analyse the cost-efficacy trade-off in a search process. We have derived $\xi$ from considering the efficacy, $P_{s}$, weighted by an exponential function of the cost expectation value $\langle C\rangle$. The simple framework of diffusion under stochastic resetting with rate $r$ has allowed us to derive analytical expressions for $P_{s}$ and $\langle C\rangle$ and thus to carry out the minimization of $\xi$. This has resulted in surprisingly rich phase diagrams, exhibiting classical continuous and discontinuous transitions, but also non-standard transitions with re-entrant behaviour and overhangs. These transitions imply changes of the optimal search strategy that may be gradual or sudden. We have shown that these transitions may be understood within a simple Landau-like expansion of the profligacy. As the profligacy just requires a cost and a measure of success as inputs, it has the potential for application in wider contexts. It would be of interest to see if the different classes of transition that we have identified here, arise more generally.

### v2

@@ VI Landau-like Expansion
with all the coefficients $a_{i}=a_{i}(\lambda,M)$. Due to the absence of an $R\to-R$ symmetry, we have to include all terms of the expansion. Related Landau-like expansions have been made in [26, 31, 47]. The expansion of $\langle C\rangle$ is obtained directly from (3), (4) and (6), (7). We also require the expansion of $P_{\rm s}$ for the two resetting ensembles, obtained by expanding out the integrands of (5) and (8) in terms of $R$ and integrating term by term. The details of the expansion for $P_{S}$ are provided in Appendix C.
We now discuss the different cases that occur for transitions in the global minimum of the profligacy $\xi$ as we vary $\lambda$.
VI.1 Classical Continuous Transition
@@
The simplest case is when $a_{2}>0$ in the expansion (9) and we may ignore higher order terms. The curve of continuous transitions, $\lambda^{*}(M)$, is obtained by solving $a_{1}(\lambda,M)=0$ for $\lambda$, as for a continuous transition seen in equilibrium systems [48]. This simple scenario pertains in the case of RBM with a linear cost, where we saw in Fig. 2 that a continuous transition occurs on increasing $\lambda$, for sufficiently low $M$.
Figure 4: The figure presents classical continuous transition observed in $\xi$ for the case of RBB with linear cost per reset when $\lambda$ is varied for $M=1.0$. (a) The global minimum which is at $R^{*}=0$ initially continuously transitions to the new global minimum $R^{*}$ in (b) as $\lambda$ is varied.
For the case of RBM the continuous transition condition, $a_{1}(\lambda,M)=0$, is satisfied at $\lambda^{*}(M)$ given by
$\lambda^{*}=\frac{d\expectationvalue{C}^{\text{RBM}}}{dR}\bigg{|}_{R\to 0^{+}}\times\frac{\text{erfc}\left(\frac{M}{\sqrt{2}}\right)}{g_{1}(M)}\;,$ (10)
with
$\frac{d\expectationvalue{C}^{\text{RBM}}_{\text{lin}}}{dR}\bigg{|}_{R\to 0^{+}}=\frac{4}{3\sqrt{\pi}}\;,$ (11)
and
$\frac{d\expectationvalue{C}^{\text{RBM}}_{\text{quad}}}{dR}\bigg{|}_{R\to 0^{+}}=1\;.$ (12)
However, since we consider the investment $\lambda$, to act as a penalty if the searcher does not find the target, we require $\lambda>0$. But $g_{1}(M)<0$ for $M>0.8198\dots$ which suggests that a continuous transition cannot exist for RBM if the rescaled distance to the target is greater than the threshold value $M_{T}=0.8198\dots$. Further, $a_{1}(\lambda^{*},M)>0$ for all values $M<M_{T}$, which implies that the system has a continuous transition for all values of $M<M_{T}$. At the threshold value, $\lambda^{*}(M_{\rm T})\to\infty$ and for $M>M_{T}$ no transition occurs. The exact value $M_{\rm T}=0.8198\ldots$ that is obtained from this procedure provides the horizontal line in Fig. 2(a). This exactly matches with the result obtained in the FIG. 2(a) and 2(b) where no transitions are observed beyond the threshold value, $M_{T}$.
For the case of RBB, the continuous transition condition, $a_{1}(\lambda,M)=0$, is satisfied at $\lambda^{*}(M)$. There we obtain the curve of continuous transition ($a_{1}(\lambda,M)=0$) to be
$\lambda^{*}=\frac{d\expectationvalue{C}}{dR}\bigg{|}_{R\to 0^{+}}\times\frac{e^{-2M^{2}}}{h_{1}(M)}\;,$ (13)
where we have the derivatives
$\frac{d\expectationvalue{C}^{\text{RBB}}_{\text{lin}}}{dR}\bigg{|}_{R\to 0^{+}}=\sqrt{\pi}\;,$ (14)
and
$\frac{d\expectationvalue{C}^{\text{RBB}}_{\text{quad}}}{dR}\bigg{|}_{R\to 0^{+}}=\frac{8}{3}\;.$ (15)
Since $h_{1}(M)>0$ for all values of $M$, we obtain a solution for $a_{1}=0$ for all values of $M$ for RBB. For small values of $\lambda^{*}$ this corresponds to a continuous transition, however for large values of $\lambda^{*}$ we find that $a_{2}<0$ and we have to consider a classical discontinuous transition which we now discuss.
VI.2 Classical Discontinuous Transition
@@
The next case is where $a_{2}$ may be positive or negative according to parameters, but $a_{3}$ is positive. This is the case for RBB with linear cost, Fig. 2(c). For $a_{2}>0$ we obtain a continuous transition at $a_{1}=0$ but for $a_{2}<0$ there is a discontinuous transition.We then have a local minimum in the profligacy at a non-zero value of $R$ in addition to a boundary minimum at $R=0$. If on increasing $\lambda$ the global minimum switches between these two local minima, we have a jump in $R^{*}$ from a zero to a non-zero value. This is the classical scenario for a discontinuous transition. A classical tricritical point occurs when a continuous transition line meets a discontinuous transition line, which occurs when $a_{1}=a_{2}=0$.
Figure 5: The figure presents classical discontinuous transition observed in $\xi$ for the case of RBB with linear cost per reset when $\lambda$ is varied for $M=0.63$. (a) The global minimum which is at $R^{*}=0$ initially discontinuously transitions to the new global minimum $R^{*}$ in (b) when the local minima becomes the new global minima as $\lambda$ is varied.
A classical tricritical point occurs in the case of RBB with linear cost as is seen in Fig. 2(c). Then for $a_{2}>0$ we obtain a continuous transition at $a_{1}=0$ but for $a_{2}<0$ there is a discontinuous transition and if $a_{1}=a_{2}=0$ (and $a_{3}>0$), we obtain a tricritical point. To obtain the tricritical point for RBB with linear cost per reset, we set $a_{2}(\lambda^{*},M)=0$, which gives
$e^{-2M^{*2}}\frac{d^{2}\expectationvalue{C}}{dR^{2}}\bigg{|}_{R\to 0^{+}}+e^{2M^{*2}}\lambda^{*}h_{1}^{2}(M^{*})-\lambda^{*}h_{2}(M^{*})=0\;,$ (16)
which upon solving yields the tricritical point $(M^{*},\lambda^{*})=(0.6333,8.4994)$ (filled circle in Fig. 2(c)). For $\lambda$ approaching this point from above we have a classical discontinuous transition
VI.3 Non-classical Discontinuous Transition
@@
The interesting non-classical overhangs that occur for both RBM and RBB with quadratic cost per reset (Figs. 2(b) and 2(d)) can be attributed to the coefficients satisfying $a_{3}<0$ and $a_{4}>0$. If we look for a tricritical point by solving $a_{2}(\lambda^{*},M)=0$ for $M$, we find at all such $M^{*}$ that $a_{3}(\lambda^{*},M^{*})<0$. This violates the condition for a tricritical point, which is why overhangs emerge instead.
This non-standard scenario occurs when $a_{3}<0$ and $a_{4}>0$ and we retain terms up to $R^{4}$ in the Landau-like expansion. This is the case for both RBM and RBB with quadratic cost per reset in an intermediate range of $M$. The quartic expansion in $R$ allows two local minima in the profligacy if $a_{1}<0$, or one local minimum, plus a boundary minimum at $R=0$, if $a_{1}>0$. In the latter case either a discontinuous transition to a non-zero value of $R^{*}$ can occur when the local minimum becomes the global minimum, or a continuous transition can occur when $a_{1}$ becomes negative; in the former case a discontinuous transition between two non-zero values of $R^{*}$ can occur when the global minimum switches between the two local minima. A sequence of a continuous transition (when $a_{1}$ turns negative) followed by a discontinuous transition (when the global minimum switches between the subsequent two local minima) generates the overhangs seen in Figs. 2(b),2(d). Thus the red dots in Figs. 2(b),2(d) are not classical tricritical points as can be verified by evaluating $a_{2}(\lambda^{*},M)=0$. For RBM evaluating this results in two different values $M^{*}_{1,2}$, but evaluating the coefficient of $R^{3}$ gives $a_{3}(\lambda^{*},M^{*}_{1,2})<0$.
Figure 6: The figure presents the non-classical discontinuous transition occurs observed in RBB with quadratic cost per reset as $\lambda$ is varied for $M=0.707$. (a) The global minimum initially emerges continuously from $R=0$ as $\lambda$ is varied. (b) As lambda is further increased, the second minimum becomes the global minima and there is a discontinuous transition of the optimal value $R^{*}$ to the new global minimum.
The behaviours that we have obtained within the Landau-like expansion and with linear and quadratic cost functions are generic. That is, we have explored all the possible behaviours coming from the different values $a_{i}$’s could take, for $i=1\ldots 3$ with $a_{4}$ positive. To obtain different phenomenology would require cost functions where higher coefficients have to be taken into account, and possibly more exotic transitions could be obtained. Of course the predictions of a Landau-like theory for discontinuous transitions will only be quantitatively accurate for small $R$, nevertheless the correct qualitative behaviour is predicted.
VII Summary and Outlook
@@
In summary, we have introduced the profligacy $\xi$ (1) as a tool to analyse the cost-efficacy trade-off in a search process. We have derived $\xi$ from considering the efficacy, $P_{s}$, weighted by an exponential function of the cost expectation value $\langle C\rangle$. The simple framework of diffusion under stochastic resetting with rate $r$ has allowed us to derive analytical expressions for $P_{s}$ and $\langle C\rangle$ and thus to carry out the minimization of $\xi$. This has resulted in surprisingly rich phase diagrams, exhibiting classical continuous and discontinuous transitions, but also non-standard transitions with re-entrant behaviour and overhangs. These transitions imply changes of the optimal search strategy that may be gradual or sudden. We have shown that these transitions may be understood within a simple Landau-like expansion of the profligacy.
As the profligacy just requires a cost and a measure of success as inputs, it has the potential for application in wider contexts. So far we have considered resetting Brownian motion and resetting Brownian bridge searches where we are able to compute the profligacy analytically. The Landau-like theory implies that the observed transitions should be generic. It would be of interest to see if the different classes of transition that we have identified here, arise more generally for other searches and cost functions.
@@ Acknowledgements.
Appendix A Derivation of the Mean Cost
The key quantity that is required to determine the mean total cost incurred in a diffusive resetting process is $\Omega_{r}(C,t_{f})$, the statistical weight of trajectories that start at the origin at $t=0$, return to the origin as a Poisson process at rate $r$, end at a predetermined time $t=t_{f}$ and incur a total cost $C$. These statistical weights will differ in the ensemble where the endpoint of the trajectory is free (as in Resetting Brownian Motion, RBM) or constrained to lie at the origin (as in a Resetting Brownian Bridge, RBB). Once this quantity is known, we can determine the mean cost over such trajectories as
$\expectationvalue{C}=\frac{\int{\rm d}C\,C\,\Omega_{r}(C,t_{f})}{\int{\rm d}C\,\Omega_{r}(C,t_{f})}\;.$ (17)
An explicit expression for $\Omega_{r}(C,t_{f})$ is obtained from a renewal equation. The idea is to consider the evolution from the start of the process until either one of two things happens. The first possibility is that the particle resets for the first time at some time $0\leq t\leq t_{f}$, after which the entire process restarts from the origin, with the remainder of the trajectory lasting a time $t-t_{f}$ and incurring a cost $C-c(x)$, where $c(x)\geq 0$ is the cost of resetting from the point $x$ to the origin. In this case, the particle resets with probability ${\rm e}^{-rt}r{\rm d}t$ in the interval $[t,t+{\rm d}t]$, and is distributed over space as $G(x;t)$ which is the Green function for diffusion,
$G(x;t)=\frac{1}{\sqrt{4\pi Dt}}e^{-\frac{(x-x_{0})^{2}}{4Dt}}$ (18)
where $D$ is the diffusion constant. The second possibility is the particle reaches the point $x$ at time $t_{f}$ without resetting. This event arises with probability ${\rm e}^{-rt}$, incurs zero cost and we allow only those endpoints $x$ that fall within the set $E$. For the case of RBM, $E$ is the entire real line, whereas for RBB, $E$ is the origin.
Expressing these two possibilities as a renewal equation, we find
$\Omega_{r}(C,t_{f})=\int_{0}^{t_{f}}{\rm d}t\,r{\rm e}^{-rt}\int_{-\infty}^{\infty}{\rm d}x\,G(x;t)\Omega_{r}(C-c(x),t_{f}-t)+{\rm e}^{-rt}\delta(C)\int_{E}{\rm d}x\,G(x;t_{f})\;.$ (19)
The first term is a convolution and thus the recursion can be solved by introducing the double Laplace transform
$\widetilde{\Omega}_{r}(p,s)=\int_{0}^{\infty}{\rm d}C\,{\rm e}^{-pC}\int_{0}^{\infty}{\rm d}t_{f}\,e^{-st_{f}}\Omega_{r}(C,t_{f})\;.$ (20)
Note on notation
Here we have used $\widetilde{\Omega}_{r}(N,p|x_{0},s)$ to indicate a double Laplace transform. We will also use a tilde symbol to denote single Laplace transforms of a single time variable, e.g. $\widetilde{G}(x;s)$, below. The arguments of the function should make clear the number of Laplace variables. In certain places, for convenience, we will use ${\mathcal{L}}_{t\to s}$ to indicate Laplace transform to Laplace variable $s$ and ${\mathcal{L}}^{-1}_{s\to t}$ to indicate Laplace inversion to the time domain.
Laplace transforming (19) with respect to both arguments and rearranging, we find
$\widetilde{\Omega}_{r}(p,s)=\frac{\widetilde{K}(r+s)}{1-r\widetilde{W}(p,r+s)}$ (21)
where
$\widetilde{K}(s)$ $=\int_{E}{\rm d}x\,\int_{0}^{\infty}{\rm d}t_{f}\,{\rm e}^{-st_{f}}G(x;t_{f})=\int_{E}{\rm d}x\,\widetilde{G}(x;s)$ (22)
$\widetilde{W}(p,s)$ $=\int_{-\infty}^{\infty}{\rm d}x\,{\rm e}^{-pc(x)}\int_{0}^{\infty}{\rm d}t\,{\rm e}^{-st}G(x;t)$
$=\int_{-\infty}^{\infty}{\rm d}x\,{\rm e}^{-pc(x)}\widetilde{G}(x;s)$ (23)
and we have from (18) that
$\widetilde{G}(x,s)=\frac{1}{2\sqrt{Ds}}e^{-\sqrt{\frac{s}{D}}\absolutevalue{x}}\;.$ (24)
We see that the function $\widetilde{K}(s)$ is determined by the constraint placed on the endpoint of the trajectory, and that $\widetilde{W}(p,s)$ depends on the functional form of the cost. Once these functions have been determined for the cases of interest, we can obtain the mean cost as a function of $t_{f}$ from (17) via
$\expectationvalue{C}=\frac{{\mathcal{L}}^{-1}_{s\to t_{f}}\{-\partial_{p}\widetilde{\Omega}_{r}(p,s)|_{p\to 0^{+}}\}}{{\mathcal{L}}^{-1}_{s\to t_{f}}\{\widetilde{\Omega}_{r}(p,s)|_{p\to 0^{+}}\}}\;.$ (25)
Taking the limit $p\to 0^{+}$, we find that the functions to be inverted to obtain the numerator and denominator, respectively, are
$\left.-\partial_{p}\widetilde{\Omega}_{r}(p,s)\right|_{p\to 0^{+}}=$ $\frac{r(r+s)^{2}}{s^{2}}\widetilde{K}(r+s)\times$
$\int_{-\infty}^{\infty}{\rm d}x\,c(x)\widetilde{G}(x;r+s)$ (26)
$\left.\widetilde{\Omega}_{r}(p,s)\right|_{p\to 0^{+}}=$ $\frac{r+s}{s}\widetilde{K}(r+s)\;.$ (27)
A.1 Resetting Brownian Motion (RBM)
For the case of RBM, the trajectory endpoint is unconstrained, and the integral in (22) is over all $x$. For any properly normalized distribution of endpoints we then have $\widetilde{K}(s)=1/s$ and the denominator in (25) is unity. It thus remains to compute the numerator by inverting (A) for the cost function of interest. In the main text we consider power-law cost functions,
$c(x)=2^{n/2}\absolutevalue{y}^{n}\quad\mbox{where}\quad y=\frac{x}{\sqrt{2Dt_{f}}}\;,$ (28)
specifically the linear and quadratic cases, $n=1$ and $n=2$. Substituting into (A) yields
$\expectationvalue{C}_{n}^{\rm RBM}$ $=\frac{\Gamma(n+1)}{t_{f}^{n/2}}{\mathcal{L}}^{-1}_{s\to t_{f}}\left\{\frac{r}{s^{2}}\frac{1}{(r+s)^{n/2}}\right\}$
$=\frac{\Gamma(n+1)}{R^{n/2}}\frac{R\gamma(\frac{n}{2},R)-\gamma(\frac{n}{2}+1,R)}{\Gamma(\frac{n}{2})}$ (29)
where $R=rt_{f}$ is the dimensionless resetting rate, and $\gamma(s,x)$ is the lower incomplete Gamma function,
$\gamma(s,x)=\int_{0}^{x}{\rm d}u\,u^{s-1}{\rm e}^{-u}\;.$ (30)
Note that this result is obtained by recognizing the Laplace transform in (A.1) as the convolution of $t$ with $t^{n/2-1}{\rm e}^{-rt}/\Gamma(n/2)$, evaluated at $t_{f}$. Note further that $\langle C\rangle_{0}^{\rm RBM}=R$ and that for the special cases $n=1$ (‘lin’) and $n=2$ (‘quad’) considered in the main text, we have
[…]
$\expectationvalue{C}^{\rm RBM}_{\rm quad}$ $=\frac{2\left(R+{\rm e}^{-R}-1\right)}{R}\;,$ (32)
[…]
$\gamma(s+1,x)=s\gamma(s,x)-x^{s}{\rm e}^{-x}\;.$ (33)
A.2 Resetting Brownian Bridge (RBB)
For the RBB, the calculation is a little more complex due to the constraint on the trajectory endpoint. Taking $E$ to comprise just the point at the origin in (22) we now have $\widetilde{K}(s)=1/\left(2\sqrt{Ds}\right)$ and the denominator of (25) takes the form
${\mathcal{L}}^{-1}_{s\to t_{f}}\{\widetilde{\Omega}_{r}(p,s)|_{p\to 0}\}$ $=\frac{1}{2\sqrt{D}}{\mathcal{L}}^{-1}_{s\to t_{f}}\left\{\left(1+\frac{r}{s}\right)\frac{1}{\sqrt{r+s}}\right\}$
[…]
Again, the inversion can be performed by recognizing it as a convolution.
Turning now to the numerator, we find for $c(x)$ given by (28) that
[…]
$\frac{1}{2\sqrt{D}}\frac{\Gamma(n+1)}{t_{f}^{n/2}}\times{\mathcal{L}}^{-1}_{s\to t_{f}}\left\{\frac{r}{s^{2}}\frac{1}{(r+s)^{(n-1)/2}}\right\}\;.$ (35)
Comparing with (A.1) we see that
${\mathcal{L}}^{-1}_{s\to t_{f}}\{-\partial_{p}\widetilde{\Omega}_{r}(p,s)|_{p\to 0}\}=\frac{n}{2\sqrt{Dt_{f}}}\expectationvalue{C}_{n-1}^{\rm RBM}$ (36)
and by dividing by the denominator (A.2) we find
[…]
recalling that $R=rt_{f}$. For the special cases $n=1$ and $n=2$ we have
[…]
We compare the functional forms of the mean cost between the two ensembles in Fig. 7. Fig. 7(a) shows how the mean linear cost behaves with $R$, and is seen to increase indefinitely as $R\to\infty$ as $\sqrt{R}$. Corresponding plots for the mean quadratic cost are in Fig. 7(b). Unlike the case for the linear cost, the mean total quadratic cost saturates as $R\to\infty$.
Figure 7: Comparison of mean total cost $\langle C\rangle$ for (a) linear cost and (b) quadratic cost per reset for RBM and RBB. It can be seen that the costs increase monotonically for both the cases. This is consistent with the intuition that more frequent resets will incur a higher cost. Although in the case of a linear cost the mean cost increases indefinitely with $R$ we find that in the case of a quadratic cost the mean cost saturates as $R\to\infty$.
Appendix B Derivation of Success Probabilities
B.1 Resetting Brownian Motion (RBM)
For a Resetting Brownian Motion, the probability of reaching a target at position $m$ by time $t_{f}$ can be obtained from the target’s survival probability which was given as Eq. (6) in [12]. Noting that the target surviving corresponds to an unsuccessful search, we find that the success probability is given by the inverse Laplace transform
$P_{s}^{\text{RBM}}=\int_{\Gamma}\frac{{\rm d}s}{2\pi i}{\rm e}^{st_{f}}\frac{1}{s}\frac{r+s}{r+s{\rm e}^{\sqrt{\frac{r+s}{D}}m}}\;.$ (40)
where $\Gamma$ is the Bromwich contour. By introducing rescaled variables $u=st_{f},R=rt_{f}$ and $M=m/\sqrt{2Dt_{f}}$, we obtain the form presented in the main text,
$P_{s}^{\text{RBM}}=\int_{\Gamma}\frac{{\rm d}u}{2\pi i}e^{u}\frac{1}{u}\frac{R+u}{R+u{\rm e}^{M\sqrt{2}\sqrt{u+R}}}\;.$ (41)
B.2 Resetting Brownian Bridge (RBB)
The success probability for a Resetting Brownian Bridge is provided as Eq. (54) in the Supplemental Material of Ref. [16]. In the notation of the present work, this reads
$P_{s}^{\text{RBB}}=\phi(R)\int_{\Gamma}\frac{{\rm d}u}{2\pi i}{\rm e}^{u}\frac{\sqrt{u+R}}{u}\frac{R+u{\rm e}^{-M\sqrt{2}\sqrt{u+R}}}{R+u{\rm e}^{M\sqrt{2}\sqrt{u+R}}}\;,$ (42)
where
[…]
We can plot the success probabilities (41) and (42) by performing the inverse Laplace transforms numerically [46]. In Figure 8(a), we plot the two functions obtained at fixed $M=1$. For RBM, we find that the success probability decreases monotonically with $R$, whilst for RBB, the function is peaked at some nonzero resetting rate $R$. When $M$ is reduced to $1/2$, Figure 8(b), we find that both functions are peaked.
Figure 8: Comparison of probability of finding the target $(P_{\rm s})$ as a function $R$ for (a) $M=1.0$ and (b) $M=0.5$. For RBM (purple), $P_{\rm s}$ can initially be a decreasing of increasing function function depending on the value of $M$, whereas for RBB (green) $P_{\rm s}$ always increases initially with resetting, reaches a maximum then decreases for large $R$.
Appendix C Success Probability Expansions
The expansion of $P_{\rm s}$ for the two resetting ensembles can be obtained by expanding out the integrands of (5) and (8) in terms of $R$ and integrating term by term.To perform these expansions, it is helpful to define $s=u+R$, so that they become
$P_{\text{s}}^{\text{RBM}}=$ ${\rm e}^{-R}\int_{\Gamma}\frac{{\rm d}u}{2\pi i}{\rm e}^{u}\frac{1}{s-R}\frac{s}{R+(s-R){\rm e}^{M\sqrt{2s}}}\;,$ (44)
[…]
$\int_{\Gamma}\frac{ds}{2\pi i}{\rm e}^{s}\frac{\sqrt{s}}{s-R}\frac{R+(s-R){\rm e}^{-M\sqrt{2s}}}{R+(s-R)e^{M\sqrt{2s}}}\;.$ (45)
The procedure now is to expand these expressions as a power series in $R$, and invert term-by-term. Up to second order, we find for the RBM case the expansion
$P^{\text{RBM}}_{\text{s}}=$ $g_{0}(M)+g_{1}(M)R+g_{2}(M)\frac{R^{2}}{2}+\ldots\;,$ (46)
where the coefficients $g_{i}(M)$ are given by the integrals
$g_{0}(M)=$ $\int_{\Gamma}\frac{{\rm d}s}{2\pi i}{\rm e}^{s}\frac{{\rm e}^{-\sqrt{2}M\sqrt{s}}}{s}=\text{erfc}\left(\frac{M}{\sqrt{2}}\right)$ (47)
$g_{1}(M)=$ $\int_{\Gamma}\frac{{\rm d}s}{2\pi i}{\rm e}^{s}\frac{{\rm e}^{-2\sqrt{2}M\sqrt{s}}\left(-1-{\rm e}^{\sqrt{2}M\sqrt{s}}(-2+s)\right)}{s^{2}}$ (48)
$g_{2}(M)=$ $\int_{\Gamma}\frac{{\rm d}s}{2\pi i}{\rm e}^{s}\frac{{\rm e}^{-3\sqrt{2}M\sqrt{s}}\left(2+2{\rm e}^{\sqrt{2}M\sqrt{s}}(-3+s)+{\rm e}^{2\sqrt{2}M\sqrt{s}}\left(6-4s+s^{2}\right)\right)}{s^{3}}\;,$ (49)
each of which can, like $g_{0}(M)$, be expressed in a closed form in terms of error functions. Explicit expressions are provided in an accompanying Mathematica notebook uploaded to DataShare [49].
Similarly for RBB, we have
$P^{\text{RBB}}_{\text{s}}=$ $h_{0}(M)+h_{1}(M)R+h_{2}(M)\frac{R^{2}}{2}+\ldots\;,$ (50)
where
$h_{0}(M)=$ $\int_{\Gamma}\frac{{\rm d}s}{2\pi i}{\rm e}^{s}\frac{{\rm e}^{-2\sqrt{2}M\sqrt{s}}\sqrt{\pi}}{\sqrt{s}}={\rm e}^{-2M^{2}}\;,$ (51)
$h_{1}(M)=$ $\int_{\Gamma}\frac{{\rm d}s}{2\pi i}{\rm e}^{s}\frac{{\rm e}^{-3\sqrt{2}M\sqrt{s}}\sqrt{\pi}\left(-1+{\rm e}^{2\sqrt{2}M\sqrt{s}}+{\rm e}^{\sqrt{2}M\sqrt{s}}(1-2s)\right)}{s^{3/2}}\;,$ (52)
$h_{2}(M)=$ $\int_{\Gamma}\frac{{\rm d}s}{2\pi i}e^{s}\frac{2{\rm e}^{-4\sqrt{2}M\sqrt{s}}\sqrt{\pi}\left(3+6{\rm e}^{\sqrt{2}M\sqrt{s}}(-1+s)-6{\rm e}^{3\sqrt{2}M\sqrt{s}}(-1+s)+2{\rm e}^{2\sqrt{2}M\sqrt{s}}s(-3+4s)\right)}{3s^{5/2}}$ (53)
which again have closed-form expressions.
A similar process can be used to obtain the terms beyond the quadratic term. Note that in $R\to 0$ limit, (46) and (50) reduce to known results without resetting: $P^{\text{RBM}}_{\text{s}}=\text{erfc}\left(M/\sqrt{2}\right)$ [50] and $P^{\text{RBB}}_{\text{s}}={\rm e}^{-2M^{2}}$ [16].
The coefficients $a_{i}$ in (9) can be written in terms of the terms of expansion of $P_{\text{S}}$ as
$a_{0}=$ $C_{0}-\lambda\log\left(g_{0}\right)\;,$ (54)
$a_{1}=$ $C_{1}-\lambda\frac{g_{1}}{g_{0}}\;,$ (55)
$a_{2}=$ $\frac{C_{2}}{2}-\lambda\frac{\left(-g_{1}^{2}+g_{0}g_{2}\right)}{2g_{0}^{2}}\;,$ (56)
$a_{3}=$ $\frac{C_{3}}{6}-\lambda\frac{\left(2g_{1}^{3}-3g_{0}g_{1}g_{2}+g_{0}^{2}g_{3}\right)}{6g_{0}^{3}}\;,$ (57)
$a_{4}=$ $\frac{C_{4}}{24}-\lambda\frac{\left(-6g_{1}^{4}+12g_{0}g_{1}^{2}g_{2}-3g_{0}^{2}g_{2}^{2}-4g_{0}^{2}g_{1}g_{3}+g_{0}^{3}g_{4}\right)}{24g_{0}^{4}}\;,$ (58)
where $C_{n}=d^{n}\expectationvalue{C}/dR^{n}\big{|}_{R\to 0^{+}}$. The $g_{i}$ should be substituted with $h_{i}$ in the RBB case.